OpenStudy (anonymous):

Why in example 2 from lecture 11 (max and min) the equation π*((v/π)^(2/3))+(2*v)*((v/π)^(1/3)) is equal to 3*π^(-1/3)*(v^(2/3)).

OpenStudy (ybarrap):

If this relation were true, it would be so for any v; however, it's only true for \(v = \frac{(3-\pi^{\frac{2}{3}})^{\frac{3}{2}}}{2}\) or \(v = 0\). For example, if v=1 then the we must have \(\pi ^{\frac{ 2 }{ 3 }}+2=3\), which is not true.